# The Quantum Complexity of String Breaking in the Schwinger Model

### v1

@@ Abstract
String breaking, the process by which flux tubes fragment into hadronic states, is a hallmark of confinement in strongly-interacting quantum field theories. We examine a suite of quantum complexity measures using Matrix Product States to dissect the string breaking process in the 1+1D Schwinger model. We demonstrate the presence of nonlocal quantum correlations along the string that may affect fragmentation dynamics, and show that entanglement and magic offer complementary perspectives on string formation and breaking beyond conventional observables.
@@ I Introduction
The formation of flux tubes, or chromoelectric strings, between color charges is a primary emergent feature of quantum chromodynamics (QCD) [1, 2]. The nonlinearity of the Yang-Mills Lagrange density [3], specifically the gluon self interactions, combined with quantum fluctuations, confine the lines of flux between static color charges into one-dimensional “strings” at modest separations. As the separation between a static quark and antiquark increases over an intermediate region, (compared to the confinement scale), the energy in the flux tube grows approximately linearly. At a certain distance a state of two hadrons (without a flux tube between them) becomes energetically favored. Such final states correspond to dynamical quarks from the vacuum rearranging themselves into baryon number $\pm 1/3$ configurations around the static charges. This process is known as string breaking and is schematically shown in Fig. 1. In nature, string breaking occurs in high-energy collisions that produce high-multiplicity final states of strongly-interacting particles. While the dynamics of particle creation through string breaking, fragmentation and hadronization is well modeled [4, 5, 6, 7, 8], high-precision predictions in environments far from present-day experiments are challenging. Such predictions will be crucial at the future Electron-Ion Collider for determining nucleon partonic distributions, measuring gluon helicity contributions, and understanding diffractive dijet production [9, 10], as well as at the LHC for the discovery of new fundamental physics.
Figure 1: The formation of hadrons in the string breaking process. The energy density, indicated by the heatmap color, shows the linearly increasing potential with the separation $d$ of external charges (red and blue circles). A flux tube connecting the external charges develops (red tubes). At a certain separation, a pair of hadronic bound states is energetically favorable. Dynamical charges (circles with dashed border) are extracted from the vacuum to screen the external charges.
@@
Significant theoretical and numerical results exist for string breaking processes (for recent overviews, see Refs. [11, 12]). String breaking can be probed, for example, through the electric potential between static charges as a function of their separation, as demonstrated in numerous classic lattice-QCD simulations, e.g., Refs. [13, 14, 15, 16, 17, 18, 19]. More recently, advances in Hamiltonian simulation methods [20, 21, 22, 23, 24, 25, 26, 27] and quantum spin chains [28, 29, 30] have enabled access to real-time dynamics and quantum observables, such as entanglement measures. These Hamiltonian approaches have been recently extended to two spatial dimensions [31, 32, 33, 34, 35, 36, 37, 38]. In parallel, string breaking has received considerable experimental attention due to rapid progress in quantum technologies, and has been studied with quantum devices, including trapped-ion chains, Rydberg atom arrays, and superconducting qubit architectures, e.g., Refs. [39, 40, 41, 42, 43, 44, 45]. While both static and dynamical systems have been explored, deep connections with confinement and screening remain to be established.
@@
The evolution of quantum complexity, reflecting physical dynamics and emergent phenomena, is a sensitive probe of the nature of physical processes, e.g., Refs. [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In this work, we investigate the behavior of multipartite entanglement and magic (nonstabilizerness) in the ground state of quantum electrodynamics in 1+1D (Schwinger model) with static background charges, as the system evolves from a confining string to isolated bound states. The Schwinger model shares features with 3+1D QCD that make it a natural testbed to study confinement, charge screening, string breaking, and few-body bound states (hadrons). For these reasons, it has been a focal point in developing quantum simulations of quantum field theories, e.g. Refs. [66, 67, 68, 69, 70, 71]. We describe minimal simulation requirements for lattice studies of string breaking, and perform simulations to study the inherent quantum complexity (quantum correlations) in the wavefunction and reveal previously unknown structure. The measures of quantum complexity we have examined are found to change rapidly in the vicinity of string breaking, offering a complementary view of mechanisms involved in the formation of hadrons and the associated vacuum rearrangement.
II Simulating the Lattice Schwinger Model
II.1 The Hamiltonian and Gauss’s Law
@@
In 1+1D with open boundary conditions (OBCs), the distribution of fermion charges completely constrains the gauge field through Gauss’s law (see App. B for an analysis of boundary effects). Further, explicit gauge degrees of freedom are absent in axial gauge, leaving a nonlocal fermionic interaction in the Hamiltonian induced by the Coulomb potential. In the Kogut-Susskind formulation [72, 73], with staggered fermion discretization [74] and the Jordan-Wigner transformation [75, 76] to spin degrees of freedom, the lattice Hamiltonian of the Schwinger model is (for a derivation, see Ref. [73])
@@
where the lattice sites are labeled from $n=1$ to $N$, and the “link-sites” from $n=1$ to N1N-1. Here, $m_{\text{lat}}=m-\frac{g^{2}a}{8}$ [77] is the chirally-improved (bare) fermion mass, $m$ and $g$ are the bare mass and coupling respectively and $N$ is the number of staggered sites (corresponding to $N_{\text{phys}}=N/2$ physical sites), and $a$ is the (staggered) lattice spacing. [^1] We have not included constant terms in Eq. (1). The physical lattice spacing is twice the staggered spacing, $a_{\rm phys}=2a$, and the number of physical lattice sites is half the number of staggered sites, $N_{\rm phys}=N/2$. The physical length of the lattice is $L=aN=a_{\rm phys}N_{\rm phys}$. The last term specifies the nonlocal fermionic interaction, where
@@
are the electric field between sites $n,n+1$, and the charge on site $k$, respectively. By convention, we set the background field $E_{0}=0$. The flux tube connecting two static external charges is represented by the external electric field $E_{\text{ext},n}(d)$ [21]. [^2] Instead of writing the electric field in terms of $\hat{E}_{n}$ and $E_{\text{ext,n}}(d)$ one could also modify Gauss’s law as explained in [20]). A string of length $d$ centered on the middle of the lattice is implemented by applying the external field
@@
II.2 Minimal Lattice Requirements for Studies of String Breaking
@@
In lattice simulations, physical states must be well-contained within the volume of the lattice, so that boundary and finite-volume effects can be quantified and systematically removed. The Compton wavelength $l$ of excited states relevant to the physics of the process must be much smaller than the lattice size $L$. Furthermore, the pixelation (or “resolution”) of wavefunctions of low-lying states must be fine enough to resolve quantities of interest perturbatively close to their continuum values, meaning that $l/a_{\rm phys}\gg 1$. Therefore, the lattice parameters must satisfy $\frac{1}{L}\ll\Delta\ll\frac{1}{a_{\rm phys}}$ for a lattice simulation to reliably capture the properties of the vacuum. Here $\Delta\sim 1/l$ is the gap to the first excited state. [^3] As an example, for lattice QCD calculations at the physical values of the quark masses, with $m_{\pi}\sim 140~{\rm MeV}$, this condition implies that $m_{\pi}L\gg 1$ and $m_{\pi}a\ll 1$. With a lattice spacing of $a=0.1~{\rm fm}$, this gives $m_{\pi}a\sim 0.07\ll 1$, and 64 lattice sites in each spatial direction gives $m_{\pi}L\sim 4.5\gg 1$. When simulating confining systems with static background charges separated by a distance $d$ (or multiparticle states), boundary effects from the perturbations must be exponentially suppressed. This requires $L\gg d$. [^4] For a simple system of charges, the potential experienced by a static charge in a 1D lattice volume L is the sum over image charge contributions [78, 79, 80, 81, 82, 83, 84]. For PBC, $V^{\rm eff}(d)=\sum_{n=-\infty}^{+\infty}\ V(|d+nL|)$}. Twisted boundary conditions can be helpful in minimizing the effects of the boundary, see e.g., Ref. [85], particularly i-PBCs for which the leading contributions to single-particle observables vanish. A final restriction is that the distance between background charges is limited to integer multiples of the lattice spacing, i.e., d=(2n1)ad=(2n-1)a.
@@
In this work, we have selected parameters for the spectrum and interactions to satisfy these conditions. As a starting point, we have chosen to work with the same Hamiltonian parameters used in Ref. [21] of $g=0.09$, $m_{\text{lat}}=0.045$, $a=1$ and $N=220$ staggered lattice sites ($L=110$) in Eq. (1). These choices, along with the vector meson mass $M_{v}$ and the scalar meson mass $M_{s}$, are given in Table 1.555These masses are consistent with the scalar meson being a bound state of two vector mesons, with binding energy of $2M_{v}-M_{s}\sim 0.0188$. These simulation parameters give $M_{v}a_{\rm phys}\sim 0.155$ and $M_{v}L\sim 17.08$, which satisfy conditions for perturbative finite volume and lattice spacing corrections. Further, the finite volume corrections are suppressed by ${\cal O}(e^{-M_{v}L/2})\sim 2\times 10^{-4}$ for $d=L/2$. Continuum physics is identified from the $a\rightarrow 0$, $N\rightarrow\infty$ limit of the Hamiltonian in Eq. (1) with fixed $m/g$ (with fixed $m$ and $g$). To approach the continuum with this physical parameter set, we have performed simulations with $\{N,a\}=\{220,1\},\{440,\frac{1}{2}\}$ and $\{880,\frac{1}{4}\}$. These parameter choices keep the spatial extent of the lattice fixed to $L=110$. The corresponding values of $M_{v}$ and $M_{s}$ are shown in Table 1.
NN $a$ $m_{\text{lat}}$ $M_{s}$ $M_{v}$
220 1.00 0.04500 0.13647 0.07763
440 0.50 0.04551 0.13650 0.07919
880 0.25 0.04576 0.13651 0.07996
Table 1: Minimal lattice parameters for approaching the continuum limit in simulations of string breaking For $g=0.09$ and $m=0.04601$ and a given system size NN (first column) and lattice spacing $a$ (second column), the third column displays the chirally-improved mass $m_{\text{lat}}$. The fourth and fifth columns give the scalar and vector meson masses, $M_{s}$ and $M_{v}$, respectively.
III Measures of Quantum Complexity and Gauge Invariance
@@
The distribution of quantum information in a state can provide insight into its underlying structure. In gauge theories, physically-meaningful measures of entanglement and quantum complexity must be gauge invariant, which restricts their definition to operations that preserve global charges. The ground state wavefunction $|\psi\rangle$ of the Schwinger model has vanishing total electric charge, $Q=0$. However, reduced states have contributions from all charge sectors. For a bipartition of the lattice into regions $A$ and $B$, the reduced density matrix $\hat{\rho}_{A}$ has a block diagonal structure, with the blocks characterized by the charge within region $A$,
$|\psi\rangle$ $=$ $\sum_{Q,i}\ c^{(Q)}_{i}|\psi^{(Q,i)}_{A}\rangle\otimes|\psi^{(-Q,i)}_{B}\rangle\ \$,
$\hat{\rho}_{A}$ $=$ ${\rm Tr}_{B}\left[\hat{\rho}_{AB}\right]\ =\ \sum_{Q}\ p_{Q}\ \hat{\rho}^{(Q)}_{A}\$, (4)
where $\hat{\rho}_{AB}=|\psi\rangle\langle\psi|$. The block-diagonal form of $\hat{\rho}_{A}$ allows the entanglement entropy to be decomposed into contributions from the charge sectors [86, 87] (see also Refs. [88, 89, 90, 91, 92, 93, 94, 95, 96, 97] and references therein)
$S(\hat{\rho}_{A})$ $=$ $\sum_{Q}p_{Q}\ S(\hat{\rho}^{(Q)}_{A})-\sum_{Q}p_{Q}\log_{2}p_{Q}\ ,$ (5)
$I(A:B)$ $=$ $S(\hat{\rho}_{A})+S(\hat{\rho}_{B})-S(\hat{\rho}_{AB})\ ,$ (6)
using the definition in Eq. (5). The antiflatness for the same bipartition, ${\cal F}_{A}(\hat{\rho}_{A})$, which provides a lower bound to the nonlocal magic [61], [^6] Results from small-qubit systems suggest the nonlocal magic is related to its lower bound by ${\cal M}_{2}^{(NL)}=4{\cal F}$. becomes
${\cal F}_{A}(\hat{\rho}_{A})$ $=$ $\sum_{Q}p_{Q}^{2}\ {\cal F}_{A}(\hat{\rho}_{A}^{(Q)})\ ,$
[…]
The Robustness of Magic (RoM) is defined by the minimum distance to the surface defined by stabilizer states,
@@
and hence provides a measure of the difficulty for classical simulation (more specifically, $R(\hat{\rho})-1$). This is a faithful magic monotone for both pure states and mixed states. The nonlocal RoM (NL RoM) associated with subsystems $A$ and $B$, is defined by a further minimization with respect to charge-preserving local unitary transformations in each region,
The transformations are constrained to preserve charge in each region for a gauge-invariant definition of $R_{AB}^{({\rm NL})}$. For example, in the Schwinger model with the mapping to qubits as in Eq. (1), for regions $A$ and $B$ each having support on two qubits, the transformations are each parameterized by four angles
$\hat{U}({\bm{\theta}})$ $=$ $e^{-i\left(\theta_{0}\hat{Z}\hat{I}+\theta_{1}\hat{I}\hat{Z}+\theta_{2}\hat{Z}\hat{Z}+\theta_{3}(\hat{X}\hat{X}+\hat{Y}\hat{Y})\right)}\ \$, (11)
as opposed to the 15 angles for an arbitrary SU(4) transformation.
@@
for charge preserving transformations, as discussed around Eqs. (8) and (11). An upper bound for the nonlocal magic ${\cal M}_{2}$ is found by writing the bipartitioned $Q=0$ ground state wavefunction as in Eq. (4[99],
@@
IV Results Obtained with Classical Simulations
@@
We first show results obtained for classical local observables, conventionally studied in the context of string breaking. Figure 2 shows the vacuum-subtracted charge density $q_{n}=\langle\hat{Q}_{n}\rangle/a$ as a function of the static charge separation $d$. With the simulation parameters described in the previous section, a small range of $d$ captures significant changes in the structure of the ground state wavefunction. This rapid transition from strings to bound hadronic states takes place from d=4550d=45-50. Dramatic changes in the charge distribution reflect the change from a single quarkonium-type state with a string between the static charges to a molecular-type state [100] of two mesons. As the external charges separate beyond this point, the mesons become isolated and the string between them vanishes.
@@
Figure 2: The charge density $q_{n}$ through the string breaking process. Left: the vacuum-subtracted charge density obtained with simulation parameters described in the text using $N=880$ and $a=1/4$, as a function of position $n$ and external charge separation $d$. Right: cross sections of the charge density for a selection of $d$’s. The inset shows the slope of the charge density at the center of the lattice, which has a local extremum at $d=48.5$.
@@
The charge distribution between the static charges (right panel of Fig. 2) is observed to transition from linear to curved with increasing $d$, consistent with the localization of the charge due to confinement. The inset shows the profile of the charge distribution. The width of this peak provides a measure of the separations over which the ground state structure changes from string-dominated to meson-dominated. The accompanying energy-momentum tensor components, as well as the chiral condensate and electric field, are discussed in App. D. These classical quantities have been well-studied in the context of string breaking (e.g., [20, 21]) and our results agree with existing literature.
@@
One informative class of quantum complexity measures for these systems is the quantum correlations between bipartitions of the wavefunction about the center of the string. The entanglement entropy of such bipartitions has been studied previously (e.g., Ref. [22]), leading to important observations, including the reduced density matrix having the form of a thermal distribution at the point of string breaking [21]. For this half-lattice bipartition, we examine the (gauge-invariant) entanglement entropy using Eq. (5), the lower bound of the nonlocal magic using the antiflatness (Eq. (7)) and the upper bound of the nonlocal magic using Eq. (17).
@@
Figure 3: Bipartite measures of entanglement and quantum complexity in string breaking. The symmetric bipartition entanglement entropy, antiflatness and the upper bound on the nonlocal ${\cal M}_{2}$ using simulation parameters $N=880,a=0.25,g=0.09,m=0.04601$, as a function of separation between static sources in terms of physical spatial sites. These quantities all peak at $d=46.5$ (see Table 2).
@@
Figure 3 shows these quantities as a function of the separation between static charges. The previously identified peak in the entanglement entropy in the vicinity of string breaking is also observed in both the lower and upper bounds of nonlocal magic. [^8] Note that the peak in the charge slope happens at slightly larger distances compared to the peak of the entanglement entropy. This result is consistent with the peak in nonlocal magic being driven by the entanglement between the regions. [^9] Given the required size of the lattices, it is impractical to perform the minimization to determine the nonlocal magic. As a result, the bounds provide all of the accessible information about the bipartition nonlocal magic.
@@
We conclude that during the transition from a string-dominated to a meson-dominated ground state structure, the nonlocal quantum magic of the bipartition increases above its vacuum value, and returns to its vacuum value as the mesons separate. This is typical of a transition from long-range to short-range correlations. Moreover, this is consistent with the bipartition surface being (exponentially) insensitive to the presence of one or more mesons located in the bulk of the lattice, far from boundaries. This can be considered a magic barrier, analogous to those observed in the time evolution of nonequilibrium systems [101].

### v2

@@ Abstract
String breaking, the process by which flux tubes fragment into hadronic states, is a hallmark of confinement in strongly-interacting quantum field theories. A suite of quantum complexity measures is examined using Matrix Product States to characterize the string breaking process in the 1+1D Schwinger model. We demonstrate the presence of nonlocal quantum correlations along the string that may affect fragmentation dynamics, and show that entanglement and magic offer complementary perspectives on string formation and breaking beyond conventional observables.
@@
Introduction—The formation of flux tubes, or chromoelectric strings, between color charges is a primary emergent feature of quantum chromodynamics (QCD) [1, 2]. At modest separations between static quark-antiquark pairs, the nonlinearity of the gluon self interactions [3], combined with quantum fluctuations, confine the flux tubes into one-dimensional “strings”. As the separation increases, the energy stored in the flux tube grows approximately linearly until a two-hadron state becomes energetically favored. Such final states correspond to dynamical quarks from the vacuum rearranging themselves into baryon number $\pm 1/3$ configurations around the static charges. In nature, string breaking occurs in high-energy collisions that produce high-multiplicity final states of strongly-interacting particles. While the dynamics of particle creation through string breaking, fragmentation and hadronization is well modeled [4, 5, 6, 7, 8], high-precision predictions in environments far from present-day experiments are challenging. Such predictions will be crucial at the future Electron-Ion Collider for determining nucleon partonic distributions, measuring gluon helicity contributions, and understanding diffractive dijet production [9, 10], and for the discovery of new fundamental physics at the LHC.
@@
Significant theoretical and numerical results exist for string breaking processes (for recent overviews, see Refs. [11, 12]). Classical lattice QCD simulations examine the electric potential between static charges as a function of their separation [13, 14, 15, 16, 17, 18, 19]. Hamiltonian and tensor network methods complement these techniques, providing access to real-time dynamics and quantum observables [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and allowing studies in two dimensions [31, 32, 33, 34, 35, 36, 37, 38]. Recent advances in experimental quantum technologies have also enabled string breaking simulations on quantum hardware, e.g., Refs. [39, 40, 41, 42, 43, 44, 45]. While both static and dynamical systems have been explored, deep connections with confinement and screening remain to be established.
@@
Measures of quantum complexity are sensitive probes of emergent phenomena [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In this work, we investigate multipartite entanglement and magic in the ground state of the Schwinger model, a well-established testbed for quantum simulations of quantum field theories [66, 67, 68, 69, 70, 71], as the system evolves from a confining string to isolated bound states (hadrons). We describe minimal simulation requirements for lattice studies of string breaking, and perform simulations to study the inherent quantum complexity (quantum correlations) in the wavefunction, revealing previously unknown structure. Measures of quantum complexity are found to change rapidly in the vicinity of string breaking, offering a complementary view of mechanisms involved in the formation of hadrons and the associated vacuum rearrangement.
@@
The Hamiltonian and Gauss’s Law—In 1+1D with open boundary conditions (OBCs), the distribution of fermion charges completely constrains the gauge field through Gauss’s law. Further, explicit gauge degrees of freedom are absent in axial gauge, leaving a nonlocal fermionic interaction in the Hamiltonian induced by the Coulomb potential. In the Kogut-Susskind formulation [72, 73], with staggered fermion discretization [74] and the Jordan-Wigner transformation to spin degrees of freedom [75, 76], the lattice Hamiltonian of the Schwinger model is (for a derivation, see Ref. [73])
@@
where the lattice sites are labeled from $n=1$ to $N$, and the link indices range from $n=1$ to N1N-1. Here, $m_{\text{lat}}=m-\frac{g^{2}a}{8}$} [77] is the chirally-improved (bare) fermion mass, $m$ and $g$ are the bare mass and coupling respectively and $N$ is the number of staggered sites (corresponding to $N_{\text{phys}}=N/2$ physical sites), and $a$ is the (staggered) lattice spacing. [^1] We have not included constant terms in Eq. (1). The physical lattice spacing is twice the staggered spacing, $a_{\rm phys}=2a$, and the number of physical lattice sites is half the number of staggered sites, $N_{\rm phys}=N/2$. The physical length of the lattice is $L=aN=a_{\rm phys}N_{\rm phys}$. The last term specifies the nonlocal fermionic interaction, where
@@
are the electric field between sites $n,n+1$, and the charge on site $k$, respectively. By convention, we set the background field $E_{0}=0$. A flux tube connecting two static external charges is represented by the external electric field $E_{\text{ext},n}(d)$ [21]. [^2] Instead of writing the electric field in terms of $\hat{E}_{n}$ and $E_{\text{ext,n}}(d)$ one could also modify Gauss’s law as explained in [20]. A string of length $d$ centered on the middle of the lattice is implemented by applying the external field
@@
Minimal lattice requirements—In lattice simulations, physical states must be well-contained within the volume of the lattice, so that boundary effects can be quantified and systematically removed. The Compton wavelength $l$ of excited states relevant to the physics of the process must be much smaller than the lattice size $L$. Furthermore, the resolution of low-lying state wavefunctions must be fine enough to resolve quantities of interest perturbatively close to their continuum values, meaning that $l/a_{\rm phys}\gg 1$. Therefore, the lattice parameters must satisfy $\frac{1}{L}\ll\Delta\ll\frac{1}{a_{\rm phys}}$ for a lattice simulation to reliably capture the physics of the process. Here $\Delta\sim 1/l$ is the gap to the first excited state. [^3] As an example, for lattice QCD calculations at the physical values of the quark masses, with $m_{\pi}\sim 140~{\rm MeV}$, this condition implies that $m_{\pi}L\gg 1$ and $m_{\pi}a\ll 1$. With a lattice spacing of $a=0.1~{\rm fm}$, this gives $m_{\pi}a\sim 0.07\ll 1$, and 64 lattice sites in each spatial direction gives $m_{\pi}L\sim 4.5\gg 1$. When simulating confining systems with static background charges separated by a distance $d$ (or multiparticle states), boundary effects on the perturbations must be exponentially suppressed. This requires $L\gg d$. [^4] For a simple system of charges on a 1D lattice, the potential experienced by a static charge is the sum over image charge contributions [78, 79, 80, 81, 82, 83, 84]. For PBCs, $V^{\rm eff}(d)=\sum_{n=-\infty}^{+\infty}\ V(|d+nL|)$}. Twisted boundary conditions can be helpful in minimizing the effects of the boundary, see e.g., Ref. [85], particularly i-PBCs for which the leading contributions to single-particle observables vanish. Boundary effects are discussed in App. B.
In this work, we select parameters for the spectrum and interactions to satisfy these conditions. As a starting point, we chose $g=0.09$, $m_{\text{lat}}=0.045$, $a=1$ and $N=220$ staggered lattice sites ($L=110$) in Eq. (1).
@@
With these parameters, the vector meson mass is found to be $M_{v}=0.13647$, so that $M_{v}a_{\rm phys}\sim 0.2729$ and $M_{v}L\sim 15.0117$, satisfying the conditions for perturbative finite volume and lattice spacing corrections. Further, the finite volume corrections are suppressed by ${\cal O}(e^{-M_{v}L/2})\sim 2\times 10^{-4}$ for $d=L/2$. Continuum physics is identified from the $a\rightarrow 0$, $N\rightarrow\infty$ limit of the Hamiltonian in Eq. (1) with fixed $m$ and $g$. To approach the continuum with this physical parameter set, we perform simulations with $\{N,a\}=\{220,1\},\{440,\frac{1}{2}\}$ and $\{880,\frac{1}{4}\}$ so that $L=110$.
@@
Measures of quantum complexity—The distribution of quantum information in a state provides insight into its underlying structure. In gauge theories, physically-meaningful measures of entanglement and quantum complexity must be gauge invariant, which restricts their definition to operations that preserve global charges. The ground state wavefunction $|\psi\rangle$ of the Schwinger model has vanishing total electric charge, $Q=0$. However, reduced states have contributions from all charge sectors. For a bipartition of the lattice into regions $A$ and $B$, the reduced density matrix $\hat{\rho}_{A}$ has a block diagonal structure, with the blocks characterized by the charge within region $A$,
@@
where $\hat{\rho}_{AB}=|\psi\rangle\langle\psi|$. The block-diagonal form of $\hat{\rho}_{A}$ allows the entanglement entropy to be decomposed into contributions from the charge sectors [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]
@@
[…]
using the definition in Eq. (5). The antiflatness for the same bipartition, ${\cal F}(\hat{\rho}_{A})$, which provides a lower bound to the nonlocal magic [61], [^5] Results from small-qubit systems suggest the nonlocal magic is related to its lower bound by ${\cal M}_{2}^{(\text{NL})}=4{\cal F}$. becomes
[…]
@@
for charge-preserving transformations to maintain gauge invariance. An upper bound for the nonlocal magic ${\cal M}_{2}$ is found by writing the bipartitioned $Q=0$ ground state wavefunction as in Eq. (4[99],
@@
The Robustness of Magic (RoM) [57, 100, 101] is defined by the minimum distance to the surface of stabilizer states,
[…]
and hence provides a measure of the difficulty for classical simulation (more specifically, $R(\hat{\rho})-1$). This is a faithful magic monotone for both pure and mixed states. The nonlocal RoM (NL RoM) associated with subsystems $A$ and $B$ is defined by a further minimization with respect to charge-preserving local unitary transformations in each region,
The transformations are constrained to preserve charge in each region for a gauge-invariant definition of $R_{AB}^{({\rm NL})}$. In the Schwinger model with the mapping to qubits as in Eq. (1), for regions $A$ and $B$ each having support on two qubits, the transformations are each parameterized by four angles
[…]
as opposed to the 15 angles for an arbitrary SU(4) transformation.
@@
Results—We first show results obtained for classical local observables, conventionally studied in the context of string breaking. The left panel of Fig. 1 shows the vacuum-subtracted charge density $q_{n}=\langle\hat{Q}_{n}\rangle/a$ as a function of the static charge separation $d$. With the simulation parameters described in the previous section, a small range of $d$ captures significant changes in the structure of the ground state wavefunction. Dramatic changes in the charge distribution around d=4550d=45-50 reflect the transition from a single quarkonium-type state with a string between the static charges to a molecular-type state [102] of two mesons. As the external charges separate beyond this point, the mesons become isolated and the string between them vanishes.
@@
Figure 1: The charge density $q_{n}$ through the string breaking process. Left: the vacuum-subtracted $q_{n}$ obtained with simulation parameters described in the text using $N=880$ and $a=1/4$, as a function of position $n$ and external charge separation $d$. Right: cross sections of $q_{n}$ for a selection of $d$’s. The inset shows the slope of $q_{n}$ at the center of the lattice, which has a local extremum at $d=48.5$.
@@
The charge distribution between the static charges (right panel of Fig. 1) is observed to transition from linear to curved with increasing $d$, consistent with charge localization due to confinement. The inset shows the profile of the charge distribution. The width of this peak provides a measure of the separations over which the ground state structure changes from string-dominated to meson-dominated. The accompanying energy-momentum tensor components, chiral condensate and electric field are discussed in App. D. These classical quantities have been well-studied in the context of string breaking (e.g., [20, 21]) and our results agree with existing literature.
@@
Quantum correlations across a bipartition of the system at the string center constitute an informative class of complexity measures. The entanglement entropy of such bipartitions has been studied previously (e.g., Refs. [21, 22]). For this half-lattice bipartition, we examine the (gauge-invariant) entanglement entropy using Eq. (5), the lower bound of the nonlocal magic using the antiflatness (Eq. (7)) and the upper bound of the nonlocal magic using Eq. (12).
@@
Figure 2: Bipartite measures of entanglement and quantum complexity in string breaking. The symmetric bipartition entanglement entropy, antiflatness and the upper bound on the nonlocal ${\cal M}_{2}$ using simulation parameters $N=880,a=1/4,g=0.09,m=0.04601$, as a function of separation between static sources in terms of physical spatial sites. These quantities all peak at $d=46.5$ (see Table 1).
@@
Figure 2 shows these quantities as a function of $d$. The previously identified entanglement entropy peak in the vicinity of string breaking is also observed in both the lower and upper bounds of nonlocal magic. This result is consistent with the nonlocal magic being driven by entanglement between the regions. [^6] The bounds provide all the accessible information about bipartition nonlocal magic because the required minimizations are impractical for the lattice sizes considered in this work.
@@
We conclude that the nonlocal magic increases during string breaking before returning to its vacuum value as the mesons separate. This is typical of a transition from long-range to short-range correlations. Moreover, this is consistent with the bipartition surface being (exponentially) insensitive to the presence of one or more mesons located in the bulk of the lattice, far from boundaries. This can be considered a magic barrier, analogous to those observed in the time evolution of nonequilibrium systems [103].